OpenStudy (anonymous):

Can anyone help me with these two cause im really confused? 1. Find a third-degree polynomial equation with rational coefficients that has roots -4 and 6+i A. x^3-8x^2-11x+148=0 B. x^3-8x^2-12x+37=0 C. x^3-12x^2+37x=0 D. x^3-8x^2-11 =0 2. what does Descartes's rule of signs tell you about the real roots of the polynomial? -2x^3+3x^2-5x-2=0 A. there is one positive root and either 2 or 0 negative roots B. there are either 2 or 0 positive roots and there are either 2 or 0 negative roots C. there is one positive root and one negative root. D. there are either 2 or 0 positive

OpenStudy (anonymous):

D^and one negative root

OpenStudy (amistre64):

you can either construct an equation from the given roots, or test the roots to see which option they fit into

OpenStudy (amistre64):

the rule of sign refers to a property that descartes noticed; the number of roots has a relationship to the number of sign changes in a setup

OpenStudy (amistre64):

rule of sign doesnt tell us what the roots are, it just helps us determine a possible number -2x^3 + 3x^2 - 5x - 2 = 0 | | 1 1 there are 2 sign changes as is, so there is at most 2 positive roots, and since complex roots come in pairs, there is also a chance of 0 positive roots being taken up by 2 complex roots

OpenStudy (amistre64):

letting x be a negative term, we can assess the negative roots -2(-x)^3 + 3(-x)^2 - 5(-x) - 2 = 0 2x^3 + 3x^2 + 5x - 2 = 0 | 1 there has to be 1 negative root, which cannot be removed in pairs

OpenStudy (anonymous):

so for #2 it would be B or D?

OpenStudy (amistre64):

B suggests that it would have 2 or 0 negative roots; and we can rest assured that it has 1 negative root. So B is out D has no mention of negative roots, but it retains the idea of 2 or 0 positives which is true

OpenStudy (anonymous):

so how would you do #1 im really confused o.O

OpenStudy (amistre64):

#1 can be constructed by recalling the complex roots come in conjugate pairs, so we have 3 roots stated. Then the function has to be a product of the stated roots-x for example: spose we are given the roots (or zeroes) to be: a,b,c the function can then be defined as: f(x) = (a-x)(b-x)(c-x)

OpenStudy (anonymous):

okay so how would i do the problem :P

OpenStudy (amistre64):

what are the stated roots?

OpenStudy (anonymous):

it is -4 and 6+i?

OpenStudy (amistre64):

thats only 2 of them, you need to remember what a conjugate pair is ...

OpenStudy (anonymous):

well thats the only one it shows on the problem :P

OpenStudy (amistre64):

it gives you 1 real root, and 1 complex (imaginary) root. complex roots come in conjugate pairs ... so they have actually given you all 3 roots to play with, you simply have to recall what it means to be a conjugate pair

OpenStudy (amistre64):

so far we can construct this thing as:\[f(x)=(-4-x)(6+i-x)(??-x)\] where ?? is the conjugate of 6+i

OpenStudy (amistre64):

or, we can see if the rule of signs can help us narrow down the odds to 1 negative and 2 complexes

OpenStudy (anonymous):

so would it be 6-i-x?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

okay so f(x)=(-4-x)(6+i-x)(6-i-x) then what do you do after that

OpenStudy (amistre64):

multiply it out of course ....

OpenStudy (amistre64):

we only need the constant products since all the options have different constant values

OpenStudy (amistre64):

what is:\[-4(6+i)(6-i)\]??

OpenStudy (anonymous):

do you have multiply it ?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

did i do this right -24-24i

OpenStudy (anonymous):

-24i-24i

OpenStudy (amistre64):

not quite :) conjugate pairs produce a difference of squares: (a+b)(a-b) = a^2-b^2 (6+i)(6-i) = 6^2 - i^2 = 36 - (-1) = 37 what is 4(37) ?

OpenStudy (anonymous):

148 right?

OpenStudy (amistre64):

yes, and which option has a constant of 148?

OpenStudy (anonymous):

option A

OpenStudy (amistre64):

then option A is the correct choice

OpenStudy (anonymous):

thank you for the help :)